OpenStudy (anonymous):

What is the first step in solving this system of equations by elimination? 5a + 3b = –9 2a – 5b = –16 A. Subtract the two equations. B. Add the two equations. C. Solve for a. D. Multiply the first equation by 2 and the second by –5.

OpenStudy (compassionate):

This is so vague with the wording. 5a + 3b = –9 2a – 5b = –16 You would solve it by elimination. Thus, you would eliminate b or a.

OpenStudy (anonymous):

so d?

OpenStudy (anonymous):

ah okayy

OpenStudy (ness9630):

I agree with @wickeddd

OpenStudy (compassionate):

>> 5a + 3b = –9 2a – 5b = –16 It is D.

OpenStudy (compassionate):

Ha ha.